# Nonlinear Drift Waves Beyond the Hasegawa-Mima Model

### v2

@@ II Nonlinear equation for drift waves
$\partial_{t}(e^{\Phi}-\nabla_{\perp}^{2}\Phi)+(\nabla_{\perp}\Phi\times\hat{z}.\hat{x})+\frac{L_{n}}{\rho_{s}}(\nabla_{\perp}\Phi\times\hat{z}\cdot\nabla_{\perp})\nabla_{\perp}^{2}\Phi=0$ (13)
We define the stretched coordinates $(\xi,\eta,\tau)$ following Washimi and Tanuiti [25] for $\epsilon<<1$,
@@
Using the relations Eqs. (20, 21, 22) and ((26), Eq. (12) can be expressed in the following form,
@@
In the phase speed frame of the waves, if we assume $\frac{L_{n}}{\rho_{s}}$ of the order of $\epsilon^{-3/2}$, then in order $\epsilon^{5/2}$, we obtain,
[$[\{\partial_{\tau}\Phi_{1}-U\partial_{\xi}(\frac{1}{2}\Phi_{1}^{2}+\Phi_{2})\}+\partial_{\xi}\Phi_{2}+U\partial_{\xi}\nabla_{\perp}^{2}\Phi_{1}]$ (28)
@@
where $[-U\partial_{\xi}\Phi_{1}-\partial_{\xi}\Phi_{1}]=0$ for $U=1$ and the above equation reduces to,
$[\partial_{\tau}\Phi-\partial_{\xi}(\frac{1}{2}\Phi^{2})+\partial_{\xi}\nabla_{\perp}^{2}\Phi]+\frac{L_{n}}{\rho_{s}}\{\Phi,\nabla_{\perp}^{2}\Phi\}\simeq 0$ (29)
where $\Phi_{1}=\Phi$ has been used. The vorticity term can be of the same order as the scalar nonlinear term ($\frac{1}{2}\Phi^{2}$) if $\frac{L_{n}}{\rho_{s}}\simeq\epsilon^{-3/2}$ that is $\frac{\rho_{s}}{L_{n}}\simeq\epsilon^{3/2}$. Equation (29) is the general nonlinear equation for pure drift waves in stretched coordinates $(\eta,\xi,\tau)$. If $\epsilon^{4}\frac{L_{n}}{\rho_{s}}<<\epsilon^{5/2}$, then the Eq. (29) in order $\epsilon^{5/2}$ becomes,
$[\partial_{\tau}\Phi-\partial_{\xi}(\frac{1}{2}\Phi^{2})+\partial_{\xi}\nabla_{\perp}^{2}\Phi]\simeq 0$ (30)
The solution of nonlinear Eq. (29) is not straightforward because it contains the effects of compressibility via electron scalar nonlinear term [17] and ions vorticity term which was considered by Hasegawa and Mima [14]. The KdV equation can be obtained by simply ignoring the higher order vorticity term in Eq. (29). But Hasegawa-Mima like equation can be obtained only if the nonlinear electron density term is dropped forcefully. Both these limiting cases are presented in the following sections for the sake of completeness.
III Soliton solution
If $\Phi=\Phi(y,t)$ is assumed, then the vorticity term disappears in Eq. (29) and it reduces to Eq. (5) because $\nabla_{\perp}^{2}=\partial_{\xi}^{2}$ in this case. In moving coordinate $\chi=\xi-M\tau$ where $\frac{M_{0}}{v_{De}}$ is normalized speed of the solitary structure, the solution of Eq. (5) turns out to be,
$\Phi(\chi)=-\mid\Phi_{m}\mid Sech^{2}(\frac{\chi}{W})$ (31)
@@ III Soliton solution
IV Vortex solution
If scalar nonlinearity is ignored, then Eq. (29) reduces to,
@@ IV Vortex solution
where $\mu^{2}=(C_{0}+\lambda^{2})$ and $d_{0}=C_{0}a_{0}$. Now we transform the above equation from $(X,Y)$ to polar coordinates $(r,\theta)$ with $X=r\cos\theta$, $Y=\sin\theta$, and $\theta=\tan^{-1}\frac{Y}{X}$. In polar coordinates, Eq. (36) is,
@@
where $J_{2}(s_{0}R_{0})$ is the Bessel function of order two and $K_{2}(\lambda R_{0})$ is modified Bessel function of order two. The solution of HM equation has been discussed in several research articles and books.
V Summary
A nonlinear equation for two-dimensional drift waves has been derived to demonstrate that the Hasegawa-Mima (HM) equation does not fully capture the dynamics of these waves. Specifically, the electron nonlinear term [17] must appear alongside the ion vorticity term in the HM equation. The flaw in the original derivation by Hasegawa and Mima [14] is identified: the authors overlooked the role of the Boltzmann distribution of electrons by using the simplified relation $\frac{\tilde{n}_{i}}{n_{i0}}=\Phi$ to describe the nonlinear ion density response. However, the correct expression should be $\frac{{n}_{i}}{n_{i0}}=e^{\Phi}$ which yields $\frac{\tilde{n}_{i}}{n_{i0}}=(\Phi+\frac{1}{2}\phi^{2})$. This correction introduces a scalar nonlinear term, $\frac{1}{2}\Phi^{2}$, which must appear in Eq. (13) along with the vorticity term.
Interestingly, in the drift wave’s phase velocity frame, the scalar nonlinearity from electrons and the vector nonlinearity from ions (i.e., vorticity) are inherently coupled. Drift waves propagate predominantly in a direction perpendicular to both the background density gradient $\nabla n_{0}$ and the ambient magnetic field ${\bf B}_{0}$. If the electrostatic potential depends only on a single coordinate—as in the chosen geometry $\Phi=\Phi(y)$—then the electron nonlinearity leads to the formation of density cavities [17]. In contrast, if the electron nonlinear contribution is artificially suppressed and a two-dimensional potential $\Phi=\Phi(x,y)$ is assumed, the ion vorticity generates dipolar vortices, similar to the solutions of the Hasegawa-Mima equation.
The general nonlinear equation has been derived in the frame moving with the wave phase speed, with the normalized speed set to $U=1$. From the author’s perspective, the electron nonlinearity is the dominant effect and should not be neglected in the derivation of nonlinear drift wave equation.
@@ V Summary
    A general nonlinear equation for pure drift waves has been derived, incorporating both scalar and vector nonlinearities.
@@
    The nonlinear electron density term is generally dominant over the ion vorticity nonlinearity.
@@
    The nonlinear term $\frac{1}{2}\varphi^{2}$ is missing in the original Hasegawa-Mima equation [14] because they approximated $\frac{\partial_{t}\tilde{n_{i}}}{n_{i0}}=\partial_{t}\Phi$ whereas the quasi-neutrality condition yields $\frac{\partial_{t}\tilde{n_{i}}}{n_{i0}}\simeq\partial_{t}(\Phi+\frac{1}{2}\Phi^{2})$ assuming the inertia-less electrons to follow the Bltzmann distribution.
This work is expected to be valuable for future studies of nonlinear drift waves, which have wide-ranging applications in space, astrophysical, and laboratory plasmas.
VI Data Availability Statement

### v3

@@ 3 Comments on Hasegawa-Mima Equation
4 Nonlinear Equation in Stretched Coordinates
Here we try to obtain nonlinear drift wave equation using reductive perturbation method (RPM) in stretched coordinates $(\eta,\xi,\tau)$ for $\epsilon<<1$, by defining,
@@ 4 Nonlinear Equation in Stretched Coordinates
Using the relations (25, 26, 27) and ((31), the Eq. (12) can be expressed in the following form,
@@
If we assume $\frac{L_{n}}{\rho_{s}}$ of the order of $\epsilon^{-3/2}$, then in order $\epsilon^{5/2}$, we obtain,
$[\{\partial_{\tau}\Phi_{1}-U\partial_{\xi}(\frac{1}{2}\Phi_{1}^{2}+\Phi_{2})\}+\partial_{\xi}\Phi_{2}+U\partial_{\xi}\nabla_{\perp}^{2}\Phi_{1}]$ (33)
@@
where $U\partial_{\xi}\Phi_{1}-\partial_{\xi}\Phi_{1}]=0$ for $U=1$ and the above equation reduces to,
$\partial_{\tau}\Phi-\partial_{\xi}(\frac{1}{2}\Phi^{2})+\partial_{\xi}\nabla_{\perp}^{2}\Phi]+\frac{L_{n}}{\rho_{s}}\{\Phi,\nabla_{\perp}^{2}\Phi\}\simeq 0$ (34)
while $\Phi_{1}=\Phi$ has been used. The vorticity term can be of the same order as the scalar nonlinear term ($\frac{1}{2}\Phi^{2}$) if $\frac{L_{n}}{\rho_{s}}\simeq\epsilon^{-3/2}$ that is $\frac{\rho_{s}}{L_{n}}\simeq\epsilon^{3/2}$. Equation (34) is the general nonlinear equation for pure drift waves in stretched coordinates $(\eta,\xi,\tau)$. If $\epsilon^{4}\frac{L_{n}}{\rho_{s}}<<\epsilon^{5/2}$, then the Eq. (34) in order $\epsilon^{5/2}$ becomes,
$\partial_{\tau}\Phi-\partial_{\xi}(\frac{1}{2}\Phi^{2})+\partial_{\xi}\nabla_{\perp}^{2}\Phi]\simeq 0$ (35)
The solution of nonlinear Eq. (34) is not straightforward because it contains both the effects of compressibility via electron scalar nonlinear term [26] and ion vorticity term which was considered by Hasegawa and Mima [9]. Another new equation (35) can be obtained by simply ignoring the higher order vorticity term in Eq. (34). The Hasegawa-Mima equation can be obtained if the nonlinear electron density term is dropped. Both these limiting cases are discussed in the following sections for the sake of completeness.
5 Soliton solution
The solitary structures formed by the nonlinear drift waves were first investigated by Nozaki and Tanuiti [20] and later on by Gell [7]. These authors derived the modified KdV equations for nonlinear drift waves. But they normalized the ion velocity with ion acoustic speed $c_{s}$. However, they also used the Boltzmann distribution for electrons approximating $n_{i}\simeq n_{e}\simeq(\Phi+\frac{1}{2}\Phi^{2})$. It has been pointed out that the phase speed of the drift wave is electron diamagnetic drift $v_{De}$, therefore the ion velocity should be normalized with $v_{De}$ in the derivation of KdV equation for drift waves . For details, the reference [26] may be consulted. Equation (35) can be solved to obtain a two-dimensional bounded structure by defining a coordinate $\chi=\chi(\eta,\xi,\tau)$ moving with constant speed say, M, as,
$\chi=\lambda\eta+\xi-M\tau$ (36)
where $\lambda<1$ because the wave propagates predominantly along $\xi$ direction. Then Eq. (35) in $\chi$ frame becomes,
$Md_{x}\Phi-\frac{1}{2}d_{x}\Phi^{2}+(1+\lambda^{2})d_{x}^{3}\Phi=0$ (37)
where $d_{x}=\frac{d}{dx}$. Using boundary conditions $(\Phi,d_{x}\Phi,d_{x}^{2}\Phi)\rightarrow 0$ as $\chi\rightarrow\pm\infty$, the above equation can be expressed as,
$(\Phi^{\prime})^{2}=\frac{M\Phi^{2}+\frac{1}{3}\Phi^{3}}{(1+\lambda^{2})}$ (38)
where $\Phi^{\prime}=d_{x}\Phi$. This equation admits the solution in following form with negative $\Phi$,
[…]
where $a=\frac{\sqrt{M}}{2\sqrt{(1+\lambda^{2})}}$. For illustration, the solution has been plotted in Fig. (1) for $M=0.1$ and $\lambda=0.1$.
If $\Phi=\Phi(y,t)$ is assumed, then the vorticity term disappears in Eq. (34) automatically and it reduces to Eq. (5) because $\nabla_{\perp}^{2}=\partial_{\xi}^{2}$ in this case. In moving coordinate $\chi=\xi-M\tau$ where $\frac{M_{0}}{v_{De}}$ is normalized speed of the solitary structure, the solution of Eq. (5) turns out to be,
[…]
@@ 5 Soliton solution
Figure 1: The 3D plot of negative electrostatic potential $\Phi$ vs moving coordinate $\chi$ is shown for $M=0.1$ and $\lambda=0.1$.
6 Vortex solution
In our point of view, Eq. (13) does not yield the HM Eq. (10) because the scalar nonlinear term is not negligible, in general. For the sake of completeness, we explore what happens if this term is ignored. In this case, we get from Eq. (34), an equation similar to HM equation in stretched coordinates as follows,
@@ 6 Vortex solution
where $\mu^{2}=(C_{0}+\lambda^{2})$ and $d_{0}=C_{0}a_{0}$. Now we transform the above equation from $(X,Y)$ to polar coordinates $(r,\theta)$ with $X=r\cos\theta$, $Y=\sin\theta$, and $\theta=\tan^{-1}\frac{Y}{X}$. In polar coordinates, Eq. (45) becomes,
@@
where $J_{2}(s_{0}R_{0})$ is the Bessel function of order two and $K_{2}(\lambda R_{0})$ is modified Bessel function of order two. The solution of HM like equation has been discussed in several research articles and books.
7 Summary
A geeneral nonlinear equation for two-dimensional drift waves has been derived to demonstrate that the Hasegawa-Mima (HM) equation does not fully capture the dynamics of these waves. Specifically, the electron nonlinear term $\frac{1}{2}\Phi^{2}$ must also appear along with the ion vorticity term in the nonlinear equation for drift waves. Since Hasegawa and Mima [9] normalized the density term in ion continuity equation with the total density $n_{i}=n_{0}(x)+\tilde{n_{i}}(x,y,t)$ and used $\frac{\partial_{t}n_{i}}{n_{i}}=\partial_{t}\ln n_{i}=\partial_{t}\Phi$, therefore the nonlinear electron density term was artificially killed.
In our opinion, the correct expression is $\frac{{n}_{i}}{n_{i0}}=e^{\Phi}$ which yields $\frac{\tilde{n}_{i}}{n_{i0}}\simeq(\Phi+\frac{1}{2}\Phi^{2})$. It may be mentioned here that Washim-Tanuiti [32], Saleem [26] and many other authors follow the same relation to derive nonlinear equations for low frequency electrostatic waves. This correction introduces a scalar nonlinear term, $\frac{1}{2}\Phi^{2}$, in the Hasegawa-Mima equation. Then the complete nonlinear equation for drift waves takes the form of Eq. (13).
Interestingly, in the drift wave’s phase velocity frame, the scalar nonlinearity from electrons Boltzmann distribution and the vector nonlinearity from ion vorticity are inherently coupled. If the electrostatic potential depends only on a single coordinate—as in the chosen geometry $\Phi=\Phi(y)$—then the electron nonlinearity leads to the formation of density cavities [26]. In contrast, if the electron nonlinear contribution is suppressed and a two-dimensional potential $\Phi=\Phi(x,y)$ is assumed, the ion vorticity generates dipolar vortices, similar to the solution of Hasegawa-Mima equation. Furthermore, when the ion vorticity is assumed to be a small effect, Eq. (13) gives a two dimensional equation (35) which admits solution in the form of two dimensional soliton shown in Fig. (1). From the author’s perspective, the electron nonlinearity is the dominant effect and should not be neglected in the derivation of nonlinear drift wave equation.
@@ 7 Summary
    The general nonlinear equation (13) for pure drift waves has been derived incorporating both nonlinear terms $\frac{1}{2}\partial_{t}\Phi^{2}$ and $\frac{L_{n}}{\rho_{s}}(\nabla_{\perp}\Phi\times\hat{z}\cdot\nabla_{\perp})\nabla_{\perp}^{2}\Phi$. Equation (13) reduces to Hasegawa-Mima equation (10) if the term $\partial_{t}(\frac{1}{2}\Phi^{2})$ is ignored.
@@
    The nonlinear electron density term is generally dominant over the ion vorticity term particularly for $\rho_{s}^{2}k_{\perp}^{2}<1$.
@@
    The nonlinear term $\frac{1}{2}\partial_{t}\Phi^{2}$ is missing in the original Hasegawa-Mima equation because they approximated $\frac{\partial_{t}\tilde{n_{i}}}{n_{i0}}=\partial_{t}\Phi$ whereas the quasi-neutrality condition yields $\frac{\partial_{t}\tilde{n_{i}}}{n_{i0}}\simeq\partial_{t}(\Phi+\frac{1}{2}\Phi^{2})$ assuming the inertia-less electrons to follow the Bltzmann distribution. The nonlinear equation (13) can yield solitons in one limit (when ion vorticity is ignored) and vortices in another limit (when nonlinear density term is ignored).
This work is expected to be very useful for future studies of nonlinear drift waves which have wide-ranging applications in space, astrophysical, and laboratory plasmas.
8 Data Availability Statement